OpenStudy (anonymous):

01.03 MA.912.A.3.1 Solve –3(2x – 1) = x + 10. You must show all work.

OpenStudy (anonymous):

i think it's –3(2x – 1) = x + 10 -3 -3 (2x - 1) = (x + 10)/3 -x -x x - 1 = 10/3 +1 +1 x = 3 1/3 ( This is 3 and one thirds )

OpenStudy (anonymous):

but it's not right, i think

OpenStudy (anonymous):

Distribute the -3 to the (2x-1) = -6x+3

OpenStudy (anonymous):

–3(2x – 1) = x + 10-- distribute -6x+3=x+10--- now combine like terms adding 6x to both sides so the 6 in the left is cancel out and ur left with 3=7x+10-- now subtract 10 to both sides so the 10 in the right equals 0 and its cancel out... ur left with -7=7x--- divide by 7 so -1=x!!!

OpenStudy (anonymous):

Yea, what she said lol

OpenStudy (anonymous):

did it help or you still don't understand?

OpenStudy (anonymous):

\[-3(2x-1)=x+10\] \[-6x+3=x+10\] \[3=7x+10\] \[-7=7x\] \[\frac{-7}{7}=x\] \[-1=x\]

OpenStudy (anonymous):

i'm good